OpenStudy (anonymous):

medal which number is not rational? 5.85 63.4 8.52624 27.5

OpenStudy (anonymous):

What is a rational number?

OpenStudy (anonymous):

Can you tell me?

OpenStudy (dangerousjesse):

A rational number has a finite set of digits... Do any of these numbers have lines over them?

OpenStudy (anonymous):

nope

OpenStudy (dangerousjesse):

Then we'll say \(\frac{x}{y}\) and \(y\neq 0\) but \(y=\textsf{an integer}\)

OpenStudy (dangerousjesse):

Which of these numbers cannot be divided by an integer to form a finite number?

OpenStudy (anonymous):

I'll be honest. I have no idea.

OpenStudy (anonymous):

Because you didnt do the work in your book you just came here for an answer

OpenStudy (dangerousjesse):

You don't know that, Mason :) Rational numbers can be confusing.

OpenStudy (anonymous):

the teacher I have is a nice guy and all but he isn't all that great of a teacher. I'm trying to learn this stuff but the two people that help me with it just happen to be gone for a bit.

OpenStudy (dangerousjesse):

Let's say \(\frac{x}{2}\) when \(x=\textsf{the choices given}\) Any irrational number will have an impossible quotient (ex, 3.1415926535897932384626433832795028841971693993751 or \(\pi\))

OpenStudy (dangerousjesse):

Divide all of these numbers by two to see which one isirrational.

OpenStudy (anonymous):

8.52624 is the irrational one then?

OpenStudy (dangerousjesse):

Yes :)

OpenStudy (anonymous):

Thank you very much :D